OpenStudy (darkbeautystar):

1. Mike want to build a rectangular garden box with a total area of 250 square feet. If the length is 25 feet what must be the width?

OpenStudy (darkbeautystar):

@TheSmartOne--@mathmath333-- would you guys please help me??

OpenStudy (darkbeautystar):

@mathmath333

OpenStudy (darkbeautystar):

Phi you left..and i need help

OpenStudy (anonymous):

The answer is ten. because 10 * 25 is 250

OpenStudy (darkbeautystar):

I need to know the last sentances answer

OpenStudy (phi):

did you not follow ? first write down the formula for area of a rectangle, like this: width * length = area then replace the length with 25 and the area with 250 (you get those numbers from the problem) , like this: 25 W = 250 then divide both sides by 25 to get W= 10

OpenStudy (darkbeautystar):

Ok thanksyou--I have 2 more questions

OpenStudy (anonymous):

Go ahead

OpenStudy (darkbeautystar):

a)Write an equation that shows the relation between the relation of the numbers using w as the variable for width.

OpenStudy (phi):

we already did that. start with the formula for area of a rectangle L*W = A then replace the letters with numbers (for the ones we know) we know L is 25 and A is 250, so we write 25W= 250 that is what they want

OpenStudy (darkbeautystar):

oops sorry . didn't mean to puit that

OpenStudy (darkbeautystar):

Write an equation for any rectangle using a for area, l for length and w for width

OpenStudy (darkbeautystar):

@phi--you keep leaving

OpenStudy (darkbeautystar):

Oh wait not that q

OpenStudy (darkbeautystar):

d) solve for length in a garden box where the area is 100 square feet and the width is 12.5 feet

OpenStudy (darkbeautystar):

@phi-@phi-@phi

OpenStudy (darkbeautystar):

This isn't that hard

OpenStudy (phi):

can you do the same thing: write down the formula for the area of a rectangle then put in the numbers. can you do that ?

OpenStudy (darkbeautystar):

L*W=A L*12.5=100 sqft

OpenStudy (phi):

now divide both sides by 12.5

OpenStudy (phi):

can you do that ?

OpenStudy (darkbeautystar):

L/12.5*12.5/12.5=100

OpenStudy (phi):

you divide each side by 12.5 you start with 12.5L = 100 then write a line under each side, and write 12.5 like this \[ \frac{12.5L}{12.5} = \frac{100}{12.5} \]

OpenStudy (phi):

next, simplify the 12.5/12.5 and simplify 100/12.5 use a calculator

OpenStudy (darkbeautystar):

Is that the answer

OpenStudy (phi):

after you simplify , you get the answer

OpenStudy (darkbeautystar):

ok

OpenStudy (darkbeautystar):

while i do that can you right out this one

OpenStudy (darkbeautystar):

2. In physics, we use the equation s= d/t to show the speed of an object. This equations means that the speed of an object is equal to distance it moved divided by the time it moved. If a car was going 65 miles per hour and went 910 miles; how long was it driving?

OpenStudy (phi):

for question 2. write down the formula (that they give you) put in the numbers for the letters can you do that ?

OpenStudy (darkbeautystar):

Hangon

OpenStudy (darkbeautystar):

S=d/t 910=65/t??

OpenStudy (darkbeautystar):

@phi

OpenStudy (darkbeautystar):

@phi

OpenStudy (phi):

yes, that looks good. 910 = 65/t to find "t" we multiply both sides by t can you do that ?

OpenStudy (phi):

you do this \[ 910 t = \frac{65t}{t} \]

OpenStudy (darkbeautystar):

Is that the answer

OpenStudy (phi):

on the right side you have a t/t and that is 1 because anything divided by itself is 1 you get \[ 910 t = 65\] do you know what to do next ?

OpenStudy (darkbeautystar):

No

OpenStudy (darkbeautystar):

@phi

OpenStudy (darkbeautystar):

No I don't

OpenStudy (darkbeautystar):

Wait.... you divide both sides by 65??

OpenStudy (darkbeautystar):

@phi are you there

OpenStudy (phi):

let's start over s= d/t s is short for speed and speed is 65 d is short for distance and distance is 910 65= 910/t now multiply both sides by t you get 65 t = 910 now divide both sides by 65 t= 910/65

OpenStudy (darkbeautystar):

ok thanks---one more question

OpenStudy (darkbeautystar):

Last one!!! 3. Misha wants to know how many pepperonis to put on his pizza. He was told that he should use 1 pepperoni rounded to the nearest whole pepperoni for each .75 inches of diameter in a pizza. Show an equation that would solve how many pepperonis he would need if his pizza has a 8 inch radius.

OpenStudy (darkbeautystar):

@phi--last one!!

OpenStudy (darkbeautystar):

@mathmath333 - can you help me with bthis last q

OpenStudy (darkbeautystar):

@Preetha

OpenStudy (darkbeautystar):

Would you help with the last question